MICROWAVE SPECTRUM AND CONFORMATIONAL COMPOSITION OF ALLENYLPHOSPHINE, H_{2}C=C=CHPH_{2}

\underline{GEORGES WLODARCZAK}, DENIS PETITPREZ, JEAN DEMAISON, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Bât. P5, Université de Lille 1, FR-59655, Villeneuve d’Ascq, France ; JEAN-CLAUDE GUILLEMIN , Laboratoire de Synthèse et Activation de Biomolécules, UMR CNRS 6052, ENSCR, FR-35700 Rennes, France ; HARALD MØLLENDAL , Department of Chemistry, University of Oslo, P. O. Box 1033 Blindern, NO-0315 Oslo, Norway.

The microwave spectrum allenylphosphine has been investigated in the 8-62 GHz spectral region. Two conformers exist for this compound. The lone electron pair is syn-periplanar with the C=C-P link of atoms in one conformer, and anticlinal in the other form. A full assignment of the microwave spectrum of the more stable syn-periplanar rotamer has now been achieved. Its dipole moment has been measured. The high-K_{−1} transitions of the a-type R-branch lines of the less stable anti-clinal conformer has also been assigned. The MW work has been augmented by high-level quantum chemical calculations to determine the energies of the two conformers at the G3 level of theory. They were also useful for a first prediction of the rotational constants and the dipole moment components. Finally, the structure of both conformers has been determined.